OpenStudy (anonymous):

The current in a 100-W bulb connected to a 120-V source is __________. A. 1.2 A B. 12,000 A C. 0.5 A D. None of the options are correct.

OpenStudy (kropot72):

\[\large Power=Voltage \times Current\] \[\large Current=\frac{Power}{Voltage}=\frac{100}{120}=(you\ can\ calculate)\ A\]

OpenStudy (anonymous):

12,000A thank you !!

OpenStudy (kropot72):

Not really! What is the result when you divide 100 by 120?

OpenStudy (anonymous):

.833333 oops!!!!!!

OpenStudy (kropot72):

Good work!